OpenStudy (anonymous):

What is the equation of the line perpendicular to the line y=4/9x-2 that passes through the point(4,3)? Answers are y=-4/9x-2 y=-9/4x+6 Y=-9/4x+12 Y=-9/4x-12 Thanks

OpenStudy (anonymous):

the slope will be the inverse of the slope they give you. the inverse of 4/9 is -9/4 plug everything into this: y-y1=m(x-x1) so y-3=-9/4(x-4), then put it in y=mx+b form y-3=-9/4x+36/4; 36/4=9 y=-9/4x+12

OpenStudy (anonymous):

y1 and x1 is the point that they gave you.